# FoMo-Bench: a multi-modal, multi-scale and multi-task Forest Monitoring Benchmark for remote sensing foundation models

### v2

@@ Abstract
Forests are an essential part of Earth’s ecosystems and natural systems, as well as providing services on which humanity depends, yet they are rapidly changing as a result of land use decisions and climate change. Understanding and mitigating negative effects requires parsing data on forests at global scale from a broad array of sensory modalities, and recently many such problems have been approached using machine learning algorithms for remote sensing. To date, forest-monitoring problems have largely been addressed in isolation. Inspired by the rise of foundation models for computer vision and remote sensing, we here present the first unified Forest Monitoring Benchmark (FoMo-Bench). FoMo-Bench consists of 15 diverse datasets encompassing satellite, aerial, and inventory data, covering a variety of geographical regions, and including multispectral, red-green-blue, synthetic aperture radar (SAR) and LiDAR data with various temporal, spatial and spectral resolutions. FoMo-Bench includes multiple types of forest-monitoring tasks, spanning classification, segmentation, and object detection. To further enhance the diversity of tasks and geographies represented in FoMo-Bench, we introduce a novel global dataset, TalloS, combining satellite imagery with ground-based annotations for tree species classification, encompassing 1,000+ categories across multiple hierarchical taxonomic levels (species, genus, family). Finally, we propose FoMo-Net, a baseline foundation model with the capacity to process any combination of commonly used spectral bands in remote sensing, across diverse ground sampling distances and geographical locations worldwide. This work aims to inspire research collaborations between machine learning and forest biology researchers in exploring scalable multi-modal and multi-task models for forest monitoring. All code and data will be made publicly at https://github.com/RolnickLab/FoMo-Bench.
Keywords:
Forest monitoring Land use Remote sensing AI for Good Multi-sensor data Dataset1 Introduction
@@
Forests worldwide are under threat from climate change, land use change, and invasive species [4, 10, 24, 63], leading to significant biodiversity decline and economic losses. However, effective forest management and conservation are hampered by a lack of information, highlighting the need for scalable forest monitoring. This gap includes multiple specific problems, such as tree species identification, tree size assessment, land cover classification, etc., which also vary considerably across geographies.
Remote sensing has been explored for forest monitoring using various sensors and spatial scales [22, 43], and increasingly, deep learning methods have been used to address various specific forest-related problems [34]. Simultaneously, the field of computer vision has seen a growth in foundation models [3], i.e. models flexible to diverse modalities, scales and data regimes while performing well across downstream tasks [36, 31, 30]. However, it has been noted that general-purpose foundation models in computer vision may be ineffective in specific problem domains [55], presenting a need for foundation models explicitly designed for specific problem spaces, such as Earth observation [38] and weather or climate modeling [44].
In this work, we propose a framework to evaluate foundation models for forest-monitoring tasks. Our Forest Monitoring Benchmark (FoMo-Bench) brings together a wide range of publicly available datasets for forest monitoring, encompassing significant spatial, temporal, and spectral diversity as illustrated in Figure 1. The benchmark includes a mixture of classification, segmentation and detection tasks, from land cover classification to tree crown segmentation. To improve the geographic diversity of FoMo-Bench, we also propose TalloS, a unique, global, multi-modal dataset focused on tree species identification and based on worldwide forest inventories [32]. Each sample included in TalloS combines data acquired from the Sentinel-2 and Sentinel-1 missions, meteorological data from the ERA-5 dataset and elevation information, and georeferenced tree species information, in the form of a multi-label classification task. Finally, we use FoMo-Bench combined with unlabeled data archives to develop a baseline for multi-modal foundation models, namely FoMo-Net, which can flexibly process most commonly used spectral bands in remote sensing. Despite being constrained to a frozen encoder setting, FoMo-Net competes with supervised baselines for semantic segmentation tasks, though it underperforms supervised approaches on multi-label classification and object detection tasks. FoMo-Net acts as a baseline for an ambitious research question: Can we unify foundation models trained on specific individual remote sensors and ground sampling distances (GSD) into a singular, versatile foundation model applicable and adaptable to the requirements (e.g. combination of spectral bands, GSDs, and geographic location) of a wide range of downstream tasks? Overall, our contributions can be summarized as follows:
@@
We provide the first, curated benchmark for forest monitoring, with a global spatial coverage and a diverse set of evaluation tasks;
@@
We introduce a novel multi-modal dataset for tree species multi-label classification;
@@ 2 Related Work
Large-scale remote sensing data has been available for many years through state-funded projects such as Sentinel and Landsat. Building on such sources, various machine learning benchmarks have been developed, such as BigEarthNet [59], Sen12MS [57], Sen4AgriNet [61] and CropHarvest [66]. These initiatives have facilitated extensive research, particularly in land use and land cover (LULC) estimation and forest-related tasks such as tree species identification [14].
@@
Recent exploration of deep learning methods for forest monitoring [34] has been made possible by a range of specific datasets at various scales and geographic locations [46]. High resolution image recordings from unmanned aerial vehicles (UAVs) are part of recent datasets for automated canopy mapping [17], woody invasive species detection [35] and tree species detection [33]. In particular, ReforesTree [53] combines UAV recordings with inventories for both detecting trees and estimate their aboveground biomass. The NeonTree dataset [68] is unique in grouping airborne recordings including light detection and ranging (LiDAR), red-blue-green (RGB) and hyperspectral measurements. Data recorded from satellite sensors have been used for deforestation detection [29] and canopy height estimation [48] using Landsat, or vegetation cover estimation using Sentinel [57]. Although inventories have been used alongside satellite data in Russia with less than a thousand samples [6, 21], we are not aware of any remote sensing dataset providing fine annotations on a large volumetric and geographical scale. Recently, manual forest inventories, including manual measurements of tree height and aboveground biomass, have been aggregated in Tallo [32], a data source that we leverage for machine learning in the present work.
@@
Increasingly, benchmarks across the space of computer vision have moved beyond single tasks to encompass a diversity of related tasks. The large scale BigEarthNet-MM dataset [59], built from Sentinel-1 and Sentinel-2 data, is designed to assess both performance and training complexity of deep learning methods [47]. The SustainBench [70] combines various satellite imagery and street-level images to monitor sustainable development goals covering 15 different topics. Geo-Bench [37] proposes to group remote sensing datasets – including SAR, multispectral, elevation and cloud probabilities data – to tackle six classification and six segmentation tasks focused on LULC. However, the spatial coverage of Geo-Bench is narrow, reducing its potential for global applicability. There remains no benchmark focused on forest monitoring worldwide, combining the full suite of forest inventories, aerial and satellite data, and including the several relevant classification, segmentation and detection tasks.
Principally driven by LULC, foundation models have been investigated for multi-modal and multi-task remote sensing [60, 8] as well as focused on specific satellite products. Recent methods have exploited cropped satellite imagery to simulate various GSDs e.g. [52], as well as contrastive losses e.g. CROMA and Cross-Scale MAE [16, 62]. These methods however lack diversity in the supported spectral bands and GSDs. The recently introduced Presto algorithm [67] is able to handle a variable number of spectral bands, but is restricted to single-pixel timeseries. The SpectralGPT approach [28], leverages the diversity of multispectral bands through the use of 3D tokens, which presents scalability challenges when applied to a vast number of bands in practice. Furthermore, while SkySense [23] stands as the largest and high-performing remote sensing foundational model according to existing benchmarks to date, its architecture incorporates a distinct encoder for each group of spectral bands, which limits its adaptability to new spectral band sets introduced by novel sensors. To the best of our knowledge, foundation models have not been yet applied to multiple forest monitoring tasks at multiple resolutions (a few centimeters to tens of meters) while including a notable diversity of spectral bands, despite significant potential identified in this area [46].
@@ 3 FoMo-Bench
There is a wide variety of different open-access forest monitoring datasets, with great diversity in terms of sensors as well as spatial, spectral and temporal resolutions [46], and encompassing a variety of tasks from LULC classification to tree species identification. With the multitude of potential forest monitoring tasks, curating datasets for each is impractical. Foundation models offer the potential for a flexible backbone that can rapidly be fine-tuned on smaller annotated datasets. We here introduce FoMo-Bench, which offers a robust testbed for forest-monitoring foundation models. By carefully selecting datasets composing FoMo-Bench, we have ensured that it (1) encompasses a diversity of geographic locations and environmental conditions, (2) spans a range of task types, and (3) includes input data from a wide variety of sensors e.g. optical, SAR, multispectral, DEM and LiDAR, as well as a range of ground sampling distances, from a few centimeters to per pixel. The diversity of the FoMo-Bench data and tasks is illustrated in Figure 1 and described in more detail below. We have curated all the datasets in FoMo-Bench, spatially split them to train machine learning models and created a reproducible environment for future work with a unified set of PyTorch data loaders.
@@
Satellite-based datasets: FoMo-Bench includes seven datasets involving satellite imagery, namely BigEarthNet-MM, Sen12MS, RapidAI4EO, ForestNet, FiveBillionPixels, TreeSatAI, and TalloS. The BigEarthNet-MM [59] and Sen12MS [57] datasets are two of the most commonly used multi-modal datasets for LULC in remote sensing. Both are based on the Copernicus Sentinel missions offering SAR and multispectral data from Sentinel-1 and Sentinel-2 respectively. For the BigEarthNet-MM dataset, we adapt the existing label structure (based on the Corine land cover database [12]) to encompass 19 distinct land cover classes including broad-leaved, coniferous and mixed forests. While BigEarthNet-MM covers 10 countries in Europe, the Sen12MS dataset has a global spatial coverage (see Figure 2(a)), offering annotations for 17 classes based on MODIS land cover maps [15] with 5 different classes for forest cover. FoMo-Bench also includes RapidAI4EO [2] as a third LULC dataset with a unique source of high resolution satellite imagery offering timeseries of Planet data, at 3m resolution, for 500,000 locations across Europe. We adapt RapidAI4EO labels to encompass 24 classes, 12 of them relevant to forest monitoring. The ForestNet dataset [29] aims to identify drivers leading to deforestation and uses Landsat-8 imagery. FiveBillionPixels dataset [64] provides high resolution Gaofen-2 images () for LULC of cities in China with 24 categories including 5 forest monitoring classes. The TreeSatAI dataset [1] offers multi-modal data (e.g. aerial images, SAR and multispectral) for tree species identification, including 20 tree species derived from forest administration data in Lower Saxony, Germany. Finally, we introduce a completely novel dataset, TalloS, detailed in Section 4.
@@
Aerial-based datasets: The limited spatial resolution offered by free satellite data is not suited for certain remote sensing tasks, such as tree canopy detection or segmentation. FoMo-Bench includes eight datasets with high resolution data from aerial sensor data, namely NeonTree, Woody, ReforesTree, Spekboom, Waititu, FLAIR #1, FLAIR #2 and FORinstance. The NeonTree dataset [68] focuses on tree crown detection with RGB imagery, LiDAR and hyperspectral data recorded from airborne systems. FoMo-Bench includes the NeonTree high resolution RGB imagery (0.1m resolution) along with point cloud LiDAR data. The Woody dataset [35] is composed of RGB images recorded from an UAV, at several cm resolution, for detection of several invasive tree species in Chile. Similarly, the Spekboom [17] and Waititu [33] datasets are composed of RGB images recorded from UAVs to map tree species in South Africa and New Zealand, respectively. In the same vein, the ReforesTree dataset [53] is designed for tree crown detection with 6 classes using high-resolution RGB images recorded from an UAV, as well as ground measurements, such as tree height and diameter at breast height. The FORinstance dataset [49] provides a unique source of data with LiDAR point clouds and semantic annotations collected in 5 countries (see Fig. 2(a)). Unlike the previous aerial datasets, FLAIR #1 [20, 19] and FLAIR #2 [18, 19] are large-scale datasets testing spatiotemporal generalization. Both datasets cover wide areas in France, providing high resolution aerial images (0.2m) and elevation maps (plus Sentinel-2 timeseries in the case of FLAIR #2). These datasets test LULC classification according to 19 classes, including coniferous and deciduous forest.

### v3

# FoMo: Multi-Modal, Multi-Scale and Multi-Task Remote Sensing Foundation Models for Forest Monitoring

@@ Abstract
Forests are vital to ecosystems, supporting biodiversity and essential services, but are rapidly changing due to land use and climate change. Understanding and mitigating negative effects requires parsing data on forests at global scale from a broad array of sensory modalities, and using them in diverse forest monitoring applications. Such diversity in data and applications can be effectively addressed through the development of a large, pre-trained foundation model that serves as a versatile base for various downstream tasks. However, remote sensing modalities, which are an excellent fit for several forest management tasks, are particularly challenging considering the variation in environmental conditions, object scales, image acquisition modes, spatio-temporal resolutions, etc. With that in mind, we present the first unified Forest Monitoring Benchmark (FoMo-Bench), carefully constructed to evaluate foundation models with such flexibility. FoMo-Bench consists of 15 diverse datasets encompassing satellite, aerial, and inventory data, covering a variety of geographical regions, and including multispectral, red-green-blue, synthetic aperture radar and LiDAR data with various temporal, spatial and spectral resolutions. FoMo-Bench includes multiple types of forest-monitoring tasks, spanning classification, segmentation, and object detection. To enhance task and geographic diversity in FoMo-Bench, we introduce TalloS, a global dataset combining satellite imagery with ground-based annotations for tree species classification across 1,000+ categories and hierarchical taxonomic levels. Finally, we propose FoMo-Net, a pre-training framework to develop foundation models with the capacity to process any combination of commonly used modalities and spectral bands in remote sensing. This work aims to inspire research collaborations between machine learning and forest biology researchers in exploring scalable multi-modal and multi-task models for forest monitoring and beyond.
Code and Data — https://github.com/RolnickLab/FoMo-Bench
1 Introduction
Forests are an essential part of Earth’s ecosystems and natural systems, providing critical services on which humanity depends. They also play a vital role in biodiversity conservation, harboring countless species and maintaining ecological balance. However, forests are rapidly changing as a result of land use decisions, climate change and invasive species (Bonan 2008; Curtis et al. 2018; Hartmann et al. 2022)
@@ 1 Introduction
leading to significant biodiversity decline and economic losses. However, effective forest management and conservation are hampered by a lack of information, highlighting the need for scalable forest monitoring. This gap encompasses multiple specific problems, such as tree species identification, tree size assessment, land cover classification, etc., which also vary considerably across geographies, climatic zones and landscapes.
Remote sensing (RS) has been explored for forest monitoring using various sensors and spatial scales (Guimarães et al. 2020; Michałowska and Rapiński 2021), while increasingly, deep learning methods have been used to address various specific forest-related problems (Kattenborn et al. 2021). Simultaneously, the field of computer vision has seen a growth in foundation models (Bommasani et al. 2022), i.e. models flexible to diverse modalities, scales and data regimes while performing well across downstream tasks (Kirillov et al. 2023; Jaegle et al. 2021, 2022). However, it has been noted that general-purpose foundation models in computer vision may be ineffective in specific problem domains (Rolf et al. 2024), presenting a need for foundation models explicitly designed for specific problem spaces, such as Earth observation (Lacoste et al. 2021) and weather or climate modeling (Mukkavilli et al. 2023). Foundation models offer the potential for a flexible backbone that can rapidly be fine-tuned on smaller annotated datasets. Given the diverse nature of RS data and applications, Earth observation foundation models should be able to operate on a wide array of Ground Sampling Distances (GSD), spectral bands, image acquisition modes, and environmental conditions. Assessing the capacity of such models is far more challenging than comparing performance on standard RS datasets. In this work, we propose a framework to evaluate general remote sensing foundation models utilizing the inherent diversity of critical forest-monitoring tasks. Our Forest Monitoring Benchmark (FoMo-Bench) brings together a wide range of publicly available datasets for forest monitoring, encompassing significant spatial, temporal, and spectral diversity as illustrated in Fig. 1. FoMo-Bench includes a mixture of classification, segmentation and detection tasks, from land cover classification to tree crown segmentation. To improve the geographic diversity of FoMo-Bench, we also create TalloS, global, multi-modal dataset focused on tree species identification based on worldwide forest inventories (Jucker et al. 2022). Each sample included in TalloS combines data acquired from the Sentinel-2 and Sentinel-1 satellite missions, meteorological data from the ERA-5 (Hersbach et al. 2020) dataset, elevation information, and georeferenced tree species information for multi-label classification.
Finally, we develop FoMo-Net, a pretraining framework which can flexibly process most commonly used modalities and spectral bands in RS with a single backbone. FoMo-Net serves as an initial response to an ambitious research question: Can we unify foundation models trained on specific individual remote sensors and GSD into a singular, versatile foundation model applicable and adaptable to the requirements (e.g. combination of spectral bands, GSD, and geographic location) of a wide range of downstream tasks? To put FoMo-Net to the test, we utilize FoMo-Bench combined with unlabeled data archives to develop an initial multi-modal RS foundation model exploiting the FoMo-Net algorithm with a simple vision transformer (ViT) (Dosovitskiy et al. 2021) encoder. Following common practice in RS foundation models, we assess the capacity of our foundation model by finetuning to the FoMo-Bench tasks. While supervised baselines in FoMo-Bench set a high performance bar, FoMo-Net delivers strong predictive accuracy handling a wide array of input sensors, spectral band combinations, GSD and forest monitoring tasks. It does not require any change in the encoder, marking a crucial first step toward a multi-task, variable-modality model for forest monitoring, stimulating progress in this nascent field and laying the groundwork for future foundation models. We elaborate more on our findings and potential lines of research in Sec. 7. Overall, our contributions can be summarized as follows:
@@
We provide FoMo-Bench the first, curated benchmark for forest monitoring, with a global spatial coverage and a diverse set of evaluation tasks;
@@
We introduce TalloS, a novel multi-modal dataset for tree species multi-label classification;
@@ 2 Related Work
Large-scale RS data has been available for many years through publicly funded projects such as the Sentinel and Landsat missions. Building on such sources, various machine learning benchmarks have been developed, such as BigEarthNet-MM (Sumbul et al. 2021), Sen12MS (Schmitt et al. 2019), Sen4AgriNet (Sykas et al. 2022), and CropHarvest (Tseng et al. 2021). These initiatives have facilitated extensive research, particularly in land use and land cover (LULC) management and forest-related tasks such as tree species identification (Fassnacht et al. 2016).
@@
Forest monitoring: High resolution image recordings from unmanned aerial vehicles (UAVs) are part of recent datasets for automated canopy mapping (Galuszynski et al. 2022), woody invasive species detection (Kattenborn et al. 2019) and tree species detection (Kattenborn et al. 2020). In particular, ReforesTree (Reiersen et al. 2022) combines UAV recordings with inventories for both detecting trees and estimating their aboveground biomass. The NeonTree dataset (Weinstein et al. 2021) is unique in grouping airborne recordings including light detection and ranging (LiDAR), red-blue-green (RGB) and hyperspectral measurements. Although inventories have been used alongside satellite data in Russia with less than a thousand samples (Brieger et al. 2019), we are not aware of any RS dataset providing fine annotations on a large volumetric and geographical scale.
@@
RS benchmarks: Benchmarks across the space of computer vision have moved beyond single tasks to encompass a diversity of related tasks. The large-scale BigEarthNet-MM dataset, built from Sentinel-1 and Sentinel-2 data, is designed to assess both performance and training complexity of deep learning methods (Papoutsis et al. 2023). The SustainBench (Yeh et al. 2021) combines various satellite imagery and street-level images to monitor sustainable development goals covering 15 different topics. Geo-Bench (Lacoste et al. 2024) proposes to group RS datasets – including SAR, multispectral, elevation and cloud probabilities data – to tackle six classification and six segmentation tasks focused on LULC. However, Geo-Bench’s limited spatial coverage reduces its potential for broad global applicability. There remains no benchmark focused on forest monitoring worldwide, combining the full suite of forest inventories, aerial and satellite data, and including the several relevant classification, segmentation and detection tasks.
RS foundation models: Foundation models have been investigated for RS datasets (Sun et al. 2022; Cong et al. 2022), grounded mainly on LULC related tasks. While these approaches have shown initial success, they are often limited to specific satellite products and lack diversity in spectral bands and GSD. To address the GSD limitation, Scale-MAE (Reed et al. 2023) was trained on the FMoW-RGB dataset (Christie et al. 2018), adopting a masked image modeling (MIM) approach. This model incorporated GSD positional encoding and a Laplacian-pyramid decoder to enforce multi-scale feature learning. Similarly, Cross-Scale MAE (Tang et al. 2023) leverages FMoW-RGB to learn scale-invariant representations by integrating contrastive loss in the MIM task and introducing a cross-prediction task, where lower-resolution images predict the higher-resolution ones. SpectralGPT (Hong et al. 2024), utilizes the spectral diversity of Sentinel-2 by employing 3D tokens, which presents scalability challenges when dealing with a large number of bands. SkySense (Guo et al. 2024) stands out as the largest RS foundation model to date with 2 billion parameters. It relies on three aligned modalities i.e. WorldView RGB, Sentinel-2 multispectral and Sentinel-1 radar data, though this limits its applicability to a broader range of RS sensors. SkySense pretraining utilizes contrastive learning and cross-modal alignment. In a similar fashion, OmniSAT (Astruc et al. 2025) and CROMA (Fuller, Millard, and Green 2023) also explore contrastive and reconstruction losses to efficiently fuse information from spatially aligned RS imagery. Both models use Sentinel-1 and Sentinel-2 data, with CROMA focusing on single timestamps and OmniSAT incorporating time series and high resolution aerial data. USatMAE (Ray, Pankow, and Basu 2016) employs paired NAIP and Sentinel-2 data, using dedicated patch embeddings for each spectral band and a spectral group pooling strategy that groups bands by spatial resolution. These tokens are then processed by a transformer trained via masked image modeling. Similarly, SpectralMAE (Zhu et al. 2023) focuses on paired hyperspectral data, using a transformer encoder with masking and attention mechanisms applied exclusively to the spectral dimension. The recently introduced Presto algorithm (Tseng et al. 2023) proposes a MIM framework for paired data, with a built-in functionality to process a variable number of spectral bands, but is restricted to single-pixel time series. Finally, DOFA (Xiong et al. 2024), aims to create a sensor agnostic transformer encoder through MIM, utilizing a wavelength-conditioned dynamic patch embedding which provides flexibility regarding input spectral bands.
Recently, there has been a surge of interest in flexible yet general RS foundation models that are capable of handling diverse data sources and solving a wide range of downstream tasks. In this work, we aim to contribute to this line of research by: a) introducing an evaluation framework designed to assess the generalization ability of future models across various Earth observation modalities and conditions, with a specific focus on forest monitoring tasks that have high ecological, environmental, societal and economic impact, and b) proposing a novel pretraining paradigm that leverages any number of unaligned modalities to create sensor-agnostic foundation models. To the best of our knowledge, foundation models have yet to be trained on such a diverse range of inputs or applied to multiple downstream tasks at varying resolutions (from a few centimeters to tens of meters), while incorporating a wide array of spectral bands. This approach addresses a significant gap in the field, as highlighted by recent studies (Ouaknine et al. 2024).
@@ 3 FoMo-Bench
We introduce FoMo-Bench, which offers a robust testbed for multimodal RS foundation models, grounded on the diversity of forest-monitoring tasks. Unlike the envisioned multimodal models, conventional methods in FoMo-Bench are task-specific, delivering strong application-optimized baselines though incapable of cross-task adaptation. By carefully selecting datasets composing FoMo-Bench, we have ensured that it (1) encompasses a diversity of geographic locations and environmental conditions, (2) spans a range of task types, and (3) includes input data from a wide variety of sensors e.g. optical, SAR, multispectral, DEM and LiDAR, as well as a range of ground sampling distances, from a few centimeters to per pixel. The diversity of the FoMo-Bench data and tasks is illustrated in Fig. 1 and described in more detail below. We have curated all the datasets in FoMo-Bench, spatially split them to train machine learning models and created a reproducible environment for future work with a unified set of PyTorch data loaders. To our knowledge, this is the first benchmark grouping such diverse forest monitoring tasks both in terms of problems they address as well as in regards to the input data.
@@
Satellite-based datasets: FoMo-Bench includes seven datasets involving satellite imagery, namely BigEarthNet-MM, Sen12MS, RapidAI4EO, ForestNet, FiveBillionPixels, TreeSatAI, and TalloS. The BigEarthNet-MM and Sen12MS (Schmitt et al. 2019) datasets are two of the most commonly used multi-modal datasets for LULC in RS. Both are based on the Copernicus Sentinel missions offering SAR and multispectral data from Sentinel-1 and Sentinel-2 respectively. For the BigEarthNet-MM dataset, we adapt the existing label structure, based on the Corine land cover database (EEA 2020), to encompass 19 distinct land cover classes including broad-leaved, coniferous and mixed forests. While BigEarthNet-MM covers 10 countries in Europe, the Sen12MS dataset has a global spatial coverage (see Fig. 2(a)), offering annotations for 17 classes based on MODIS land cover maps (Friedl et al. 2010) with 5 different classes for forest cover. FoMo-Bench also includes RapidAI4EO (Bhugra et al. 2022) as a third LULC dataset with a unique source of high resolution satellite imagery offering timeseries of Planet data, at 3m resolution, for 500,000 locations across Europe. We adapt RapidAI4EO labels to encompass 24 classes, 12 of them relevant to forest monitoring. The ForestNet dataset (Irvin et al. 2020) aims to identify drivers leading to deforestation and uses Landsat-8 imagery. FiveBillionPixels dataset (Tong, Xia, and Zhu 2023) provides high resolution Gaofen-2 images () for LULC of cities in China with 24 categories including 5 forest monitoring classes. The TreeSatAI dataset (Ahlswede et al. 2023) offers multi-modal data (e.g. aerial images, SAR and multispectral) for tree species identification, including 20 tree species derived from forest administration data in Lower Saxony, Germany. Finally, we introduce a completely novel dataset, TalloS, detailed in Sec. 4.
@@
Aerial-based datasets: The limited spatial resolution offered by free satellite data is not suited for certain RS tasks, such as tree canopy detection or segmentation. FoMo-Bench includes eight datasets with high resolution data from aerial sensor data, namely NeonTree, Woody, ReforesTree, Spekboom, Waititu, FLAIR #1, FLAIR #2 and FORinstance. The NeonTree dataset (Weinstein et al. 2021) focuses on tree crown detection with RGB imagery, LiDAR and hyperspectral data recorded from airborne systems. FoMo-Bench includes the NeonTree high resolution RGB imagery (0.1m resolution) along with point cloud LiDAR data. The Woody dataset (Kattenborn et al. 2019) is composed of RGB images recorded from an UAV, at several cm resolution, for detection of several invasive tree species in Chile. Similarly, the Spekboom (Galuszynski et al. 2022) and Waititu (Kattenborn et al. 2020) datasets are composed of RGB images recorded from UAVs to map tree species in South Africa and New Zealand, respectively. In the same vein, the ReforesTree dataset (Reiersen et al. 2022) is designed for tree crown detection with 6 classes using high-resolution RGB images recorded from an UAV, as well as ground measurements, such as tree height and diameter at breast height. The FORinstance dataset (Puliti et al. 2023) provides a unique source of data with LiDAR point clouds and semantic annotations collected in 5 countries (see Fig. 2(a)). Unlike the previous aerial datasets, FLAIR #1 (Garioud et al. 2022, 2024) and FLAIR #2 (Garioud et al. 2023, 2024) are large-scale datasets testing spatiotemporal generalization. Both datasets cover wide areas in France, providing high resolution aerial images (0.2m) and elevation maps (plus Sentinel-2 timeseries in the case of FLAIR #2). These datasets test LULC classification across 19 classes.